# Zeta Zeros on the Critical Line

### v1

@@ Abstract.
Montgomery in 1973 introduced the pair correlation method to study the vertical distribution of Riemann zeta-function zeros. This work assumed the Riemann Hypothesis (RH). One striking application was a short proof that at least 2/3 of zeta-zeros are simple zeros, the first result of its type. Over the last 50 years, most work on pair correlation of zeta-zeros has continued to assume RH. Here we show that if RH could be removed from Montgomery’s simple zero proof, then this would also give a proof that 2/3 of the zeros are on the critical line. While at present we cannot unconditionally obtain this result, we have proved in joint work that this is true if we assume all the zeros are in a narrow enough region centered on the critical line. In separate joint work we use the same idea with a related pair correlation method to prove that assuming the Pair Correlation Conjecture, also without the assumption of RH, we obtain 100% of the zeros are simple and on the critical line.
@@ 1. Introduction
In 1973 Montgomery [Mon73] proved, assuming the Riemann Hypothesis (RH), that at least 2/3 of the zeros of the Riemann zeta-function are simple. Prior to this, it was not known either unconditionally or on RH that there were even infinitely many simple zeros. Currently, by a refined version of Levinson’s method, K. Pratt, N. Robles, A. Zaharescu, and D. Zeindler [PRZZ20] in 2020 proved that at least $41.7\%$ of the zeros of $\zeta(s)$ are on the critical line, and at least $40.7\%$ of the zeros are on the critical line and simple. By large scale computation with careful error analysis, Platt and Trudgian [PT21] in 2021 showed that in the critical strip up to height 3,000,175,332,800, there are exactly 12,363,153,437,138 zeros and all of them are on the critical line and are simple.
Conditionally, under the RH assumption one can obtain that at least $67.92\%$ of the zeros of $\zeta(s)$ are simple using Montgomery’s pair correlation method [CGdL20], and with a different method [CGG98, BHB13] one obtains at least $70.37\%$ of the zeros being simple. One may ask why we care about Montgomery’s pair correlation method to detect simple zeros, if one could obtain a better proportion of simple zeros using a different method. In fact, Chirre, Gonçalves, and de Laat [CGdL20]’s $67.92\%$ result appeared years after Bui and Heath-Brown [BHB13] proved the $70.37\%$ proportion, which also shows a specific interest in Montgomery’s pair correlation method. We will not dive into further details on these papers. What really has became the main motivation is the fact that Montgomery’s pair correlation method has wide applications, including direct applications to the distribution of prime numbers. This is demonstrated in several papers [GM78, HBG84, GM87, Gol88] and many more following them. In particular, Montgomery’s method has a huge flexibility that leaves a vast room for improvements, one of which we will clarify in this exposé.
In this exposé we introduce a simple method using pair correlation of zeros of $\zeta(s)$ that, on assuming additional information on the zeros, proves conditionally that there are many zeros on the critical line. Three main points to underline in our results are that 1) we can remove RH in the study of pair correlation of zeros of $\zeta(s)$, 2) we can replace RH with weaker assumptions to recover results on simple zeros that have been only obtained under RH, and that 3) for the first time, the pair correlation of zeros implies that most of the zeros lie on the critical line.
@@ 2. Zeros of the Riemann zeta-function
since this shows that $(1-2^{1-\sigma})\zeta(\sigma)>0$ so that $\zeta(\sigma)<0$ on the real line for $0<\sigma<1$. We also see immediately from the functional equation that $\zeta(0)=-1/2$. Note that this last formula gives the analytic continuation of $\zeta(s)$ in the critical strip.
If an analytic function $F(s)$ is real on a segment of the real axis it will satisfy the reflection principle that $F(\bar{s})=\overline{F(s)}$ in the symmetric region on either side of that segment where $F$ is analytical. Denoting the complex zeros of the Riemann zeta-function by $\rho=\beta+i\gamma$, we see $\zeta(\rho)=0=\zeta(\bar{\rho})$ and the functional equation gives as well $\zeta(\rho)=0=\zeta(1-\rho)$. Thus, all nontrivial zeros come in pairs symmetric with the real axis $\rho=\beta+i\gamma$ and $\bar{\rho}=\beta-i\gamma$, and zeros with $\beta\neq 1/2$ come in symmetric quadruples with respect to both the real axis and the critical line (namely, the vertical line $\sigma=1/2$): $\rho=\beta+i\gamma,\ \bar{\rho}=\beta-i\gamma,\ 1-\rho=1-\beta-i\gamma,\ 1-\bar{\rho}=1-\beta+i\gamma$. This elementary fact turns out to be fundamental later in this paper.
Since the zeros in the critical strip below the real axis are a symmetric mirror of the zeros above the real axis, we only need to study the zeros above the real axis. Thus, we let
(2.1) $N(T):=\#^{\prime}\big\{\rho=\beta+i\gamma\ :\ 0<\gamma\leq T\big\}=\sum_{0<\gamma\leq T}1,$
where $^{\prime}$ means the number of elements in the set counted with multiplicity, that is if $\rho$ is a zero with multiplicity $m_{\rho}$, then we have $m_{\rho}$ copies of $\rho$ in the sequence of zeros. Thus, for example, a triple zero gets counted 3 times in $N(T)$.
In 1905 von Mangoldt proved that, for $T\geq 3$,
(2.2) $N(T):=\sum_{0<\gamma\leq T}1=\frac{T}{2\pi}\log\frac{T}{2\pi}-\frac{T}{2\pi}+O(\log{T}),$
@@
We define the average spacing between consecutive zeros by
@@ 3. Montgomery’s RH proof of simple zeros
(3.1) $\sum_{\begin{subarray}{c}0<\gamma,\gamma^{\prime}\leq T\\ \gamma=\gamma^{\prime}\end{subarray}}1\leq\sum_{0<\gamma,\gamma^{\prime}\leq T}\left(\frac{\sin(\frac{1}{2}(\gamma-\gamma^{\prime})\log T)}{\frac{1}{2}(\gamma-\gamma^{\prime})\log T}\right)^{2}=\left(\frac{4}{3}+o(1)\right)\frac{T}{2\pi}\log T.$
Here the first inequality is trivially true. The asymptotic formula for the Fejér kernel sum over differences of zeros requires RH; the interested reader may see the proof in [Mon73]. In thinking about these double sums it may be helpful to consider the formula
(3.2) $\sum_{0<\gamma,\gamma^{\prime}\leq T}\left(\frac{\sin(\frac{1}{2}(\gamma-\gamma^{\prime})\log T)}{\frac{1}{2}(\gamma-\gamma^{\prime})\log T}\right)^{2}=2\int_{0}^{1}\bigg|\sum_{0<\gamma\leq T}T^{i\alpha\gamma}\bigg|^{2}(1-\alpha)\,d\alpha,$
@@
(3.3) $2\,{\rm Re}\int_{0}^{1}e^{iw\alpha}(1-\alpha)\,d\alpha=\left(\frac{\sin\frac{1}{2}w}{\frac{1}{2}w}\right)^{2}.$
Thus in (3.1) the double sums can be thought of as two copies of the sequence of zeros $0<\gamma\leq T$, where this sequence itself already has $m_{\gamma}$ copies of each zero of multiplicity $m_{\gamma}$.
Now consider the first sum in (3.1) with the condition that $\gamma=\gamma^{\prime}$. If we think of the terms as ordered pairs $(\gamma,\gamma^{\prime})$ where $\gamma=\gamma^{\prime}$, then a simple zero gets counted once, but a double zero with $m_{\gamma}=2$ gets counted 4 times, since $\gamma_{1}=\gamma_{2}=\gamma_{1}^{\prime}=\gamma_{2}^{\prime}$ and we get the 4 terms $(\gamma_{1},\gamma_{1}^{\prime})$, $(\gamma_{1},\gamma_{2}^{\prime})$, $(\gamma_{2},\gamma_{1}^{\prime})$, $(\gamma_{2},\gamma_{2}^{\prime})$. Thus a zero of multiplicity $m_{\gamma}$ gets counted $m_{\gamma}^{2}$ times. The proof of Montgomery’s Simple Zero Theorem is now just two lines.
@@
We have
(3.4) $\sum_{\begin{subarray}{c}0<\gamma,\gamma^{\prime}\leq T\\ \gamma=\gamma^{\prime}\end{subarray}}1=\sum_{0<\gamma\leq T}m_{\gamma}\leq\left(\frac{4}{3}+o(1)\right)\frac{T}{2\pi}\log T,$
and, using (2.1) and (2.3),
$\begin{split}\sum_{\begin{subarray}{c}0<\gamma\leq T\\ \rho\ \text{simple}\end{subarray}}1&\geq\sum_{0<\gamma\leq T}(2-m_{\gamma})=2\sum_{0<\gamma\leq T}1-\sum_{0<\gamma\leq T}m_{\gamma}\\ &\geq\left(2-\frac{4}{3}+o(1)\right)\frac{T}{2\pi}\log T=\left(\frac{2}{3}+o(1)\right)N(T)=\left(\frac{2}{3}+o(1)\right)\sum_{0<\gamma\leq T}1.\end{split}$
@@ 4. Montgomery’s Simple Zero Method if RH is not assumed
For 50 years it never seemed to occur to anyone (including the first-named author of this paper) to seriously think about this proof when RH is not assumed. As we will see in a moment, the key idea here is that, despite the title of this section, not assuming RH is not enough; what you have to do is take seriously the possibility that RH is false, and consider it without prejudice. We actually spent almost two years proving (3.1) holds with conditions weaker than RH before we realized that these results could be applied to obtain critical zeros. The source of this mental block comes from the equation
(4.1) $\sum_{\begin{subarray}{c}0<\gamma,\gamma^{\prime}\leq T\\ \gamma=\gamma^{\prime}\end{subarray}}1\ =\ \sum_{\begin{subarray}{c}\rho,\rho^{\prime}\\ 0<\gamma,\gamma^{\prime}\leq T\\ \gamma=\gamma^{\prime}\end{subarray}}1.$
The sum on the left looks like we are assuming RH because there is no $\beta$ dependence evident in the equation. The equation on the right hints at the $\beta$ dependence, since $\rho=\beta+i\gamma$ and $\rho^{\prime}=\beta^{\prime}+i\gamma^{\prime}$, but it is still easy to think RH does not matter because it is equal to the sum on the left.
We need to go back to the formulas in Section 2. With respect to the sum (4.1), we want to classify pairs of zeros which lie on the same horizontal line. Note that under the assumption of RH, these can only be the same zeros, thus
(4.2) $\underset{\gamma=\gamma^{\prime}}{\sum_{\begin{subarray}{c}\rho=\tfrac{1}{2}+i\gamma\\ 0<\gamma\leq T\end{subarray}}\sum_{\begin{subarray}{c}\rho^{\prime}=\tfrac{1}{2}+i\gamma^{\prime}\\ 0<\gamma^{\prime}\leq T\end{subarray}}}1\ =\ \sum_{\begin{subarray}{c}\rho=\rho^{\prime}\\ 0<\gamma,\gamma^{\prime}\leq T\\ \beta=\beta^{\prime}=\tfrac{1}{2}\end{subarray}}1\ =\ \sum_{\begin{subarray}{c}\rho=\tfrac{1}{2}+i\gamma\\ 0<\gamma\leq T\end{subarray}}m_{\rho}.$
Without assuming RH, there “may” be zeros $\rho=\beta+i\gamma$, $\beta\neq 1/2$, lying off the critical line, and each one of these is matched with the symmetric zero on the opposite side of the critical line $1-\overline{\rho}=1-\beta+i\gamma$. Hence, if $\beta\neq 1/2$ then the pairs of zeros $(\beta+i\gamma,1-\beta+i\gamma)$ and $(1-\beta+i\gamma,\beta+i\gamma)$ are solutions of the equation $\gamma=\gamma^{\prime}$. We call these “symmetric diagonal terms”. Thus, in addition to the sum (4.2), the summation (4.1) also includes an additional sum over these symmetric diagonal terms. Note also since we are not assuming RH, the sum over the diagonal terms is
$\sum_{\begin{subarray}{c}\rho=\rho^{\prime}\\ 0<\gamma,\gamma^{\prime}\leq T\end{subarray}}1\ =\ \sum_{\begin{subarray}{c}\rho=\beta+i\gamma\\ 0<\gamma\leq T\end{subarray}}m_{\rho}.$
Is this the end? Apparently not. Since we are working with pairs of zeros, if there are three or more zeros on the same horizontal line, we can also have terms which are not diagonal and not symmetric. To conclude, we have
(4.3) $\sum_{\begin{subarray}{c}\rho,\rho^{\prime}\\ 0<\gamma,\gamma^{\prime}\leq T\\ \gamma=\gamma^{\prime}\end{subarray}}1\ =\ \sum_{\begin{subarray}{c}\rho\\ 0<\gamma\leq T\end{subarray}}m_{\rho}+\sum_{\begin{subarray}{c}\rho\\ 0<\gamma\leq T\\ \beta\neq\frac{1}{2}\end{subarray}}m_{\rho}+\sum_{\begin{subarray}{c}\rho\neq\rho^{\prime}\\ 0<\gamma,\gamma^{\prime}\leq T\\ \ \beta+\beta^{\prime}\neq 1,\,\gamma=\gamma^{\prime}\end{subarray}}1,$
where the first sum on the right-hand side comes from the diagonal terms, the second is given by the symmetric diagonal terms, and all the rest in the third sum are non-symmetric horizontal terms.
@@
Suppose there exists a constant $C$ where $1\leq C<2$ and that, as $T\to\infty$,
(4.4) $\sum_{\begin{subarray}{c}\rho,\rho^{\prime}\\ 0<\gamma,\gamma^{\prime}\leq T\\ \gamma=\gamma^{\prime}\end{subarray}}1\leq\Big(C+o(1)\Big)\frac{T}{2\pi}\log T.$
Then asymptotically at least the proportion 2C2-C of the zeros of $\zeta(s)$ are simple; at least the proportion 2C2-C of the zeros of $\zeta(s)$ are on the critical line; and if $1\leq C<3/2$ then at least the proportion 32C3-2C of the zeros of $\zeta(s)$ are simple zeros on the critical line.
@@
We first give a remark about the restriction on the constant $C$ in our assumption. Note that
(4.5) $\sum_{\begin{subarray}{c}\rho,\rho^{\prime}\\ 0<\gamma,\gamma^{\prime}\leq T\\ \gamma=\gamma^{\prime}\end{subarray}}1~\geq~\sum_{\begin{subarray}{c}\rho,\rho^{\prime}\\ 0<\gamma,\gamma^{\prime}\leq T\\ \rho=\rho^{\prime}\end{subarray}}1~=~\sum_{\begin{subarray}{c}\rho\\ 0<\gamma\leq T\end{subarray}}m_{\rho}~\geq~\sum_{\begin{subarray}{c}\rho\\ 0<\gamma\leq T\end{subarray}}1=N(T)~\sim~\frac{T}{2\pi}\log T$
as $T\to\infty$. Thus $C$ in our assumption has to satisfy $C\geq 1$.
@@
(4.6) $\sum_{\begin{subarray}{c}\rho\\ 0<\gamma\leq T\end{subarray}}m_{\rho}+\sum_{\begin{subarray}{c}\rho\\ 0<\gamma\leq T\\ \beta\neq\frac{1}{2}\end{subarray}}m_{\rho}+\sum_{\begin{subarray}{c}\rho\neq\rho^{\prime}\\ 0<\gamma,\gamma^{\prime}\leq T\\ \ \beta+\beta^{\prime}\neq 1,\,\gamma=\gamma^{\prime}\end{subarray}}1\leq\Big(C+o(1)\Big)\frac{T}{2\pi}\log T.$
@@
$\sum_{\begin{subarray}{c}\rho\\ 0<\gamma\leq T\end{subarray}}m_{\rho}\leq\Big(C+o(1)\Big)\frac{T}{2\pi}\log T,$
which is (3.4) with $C=4/3$, and the identical argument used there gives
(4.7) $\sum_{\begin{subarray}{c}0<\gamma\leq T\\ \rho\ \text{simple}\end{subarray}}1\geq\Big(2-C+o(1)\Big)\sum_{0<\gamma\leq T}1,$
which proves that the proportion of simple zeros of $\zeta(s)$ is $\geq 2-C+o(1)$.
Next, for the symmetric diagonal terms, we have
$\sum_{\begin{subarray}{c}\rho\\ 0<\gamma\leq T\\ \beta\neq 1/2\end{subarray}}m_{\rho}\leq\Big(C+o(1)\Big)\frac{T}{2\pi}\log T-\sum_{\begin{subarray}{c}\rho\\ 0<\gamma\leq T\end{subarray}}m_{\rho}.$
Recalling (4.5) and plugging in the last inequality there into the above inequality, we obtain
(4.8) $\sum_{\begin{subarray}{c}\rho\\ 0<\gamma\leq T\\ \beta\neq 1/2\end{subarray}}m_{\rho}\leq\Big(C-1+o(1)\Big)\frac{T}{2\pi}\log T.$
@@
$\sum_{\begin{subarray}{c}\rho\\ 0<\gamma\leq T\\ \beta=1/2\end{subarray}}1=\sum_{\begin{subarray}{c}\rho\\ 0<\gamma\leq T\end{subarray}}1-\sum_{\begin{subarray}{c}\rho\\ 0<\gamma\leq T\\ \beta\neq 1/2\end{subarray}}1\geq N(T)-\sum_{\begin{subarray}{c}\rho\\ 0<\gamma\leq T\\ \beta\neq 1/2\end{subarray}}m_{\rho}\geq\Big(2-C+o(1)\Big)\frac{T}{2\pi}\log T,$
@@
and we have proved that the proportion of zeros of $\zeta(s)$ on the critical line is $\geq 2-C+o(1)$. Finally, using (4.7) and (4.8) we obtain
$\sum_{\begin{subarray}{c}\rho\\ 0<\gamma\leq T\\ \beta=\frac{1}{2}\ \text{and}\ \rho\text{ \rm simple}\end{subarray}}1~=~\sum_{\begin{subarray}{c}\rho\\ 0<\gamma\leq T\\ \rho\text{ \rm simple}\end{subarray}}1\ -\sum_{\begin{subarray}{c}\rho\\ 0<\gamma\leq T\\ \beta\neq\frac{1}{2}\ \text{and}\ \rho\text{ \rm simple}\end{subarray}}1$ $\geq~\Big(2-C+o(1)\Big)\frac{T}{2\pi}\log T-\sum_{\begin{subarray}{c}\rho\\ 0<\gamma\leq T\\ \beta\neq\frac{1}{2}\end{subarray}}m_{\rho}$
$\geq~\Big(3-2C+o(1)\Big)\frac{T}{2\pi}\log T,$

### v2

@@ Abstract.
Montgomery in 1973 introduced the pair correlation method to study the vertical distribution of Riemann zeta-function zeros. This work assumed the Riemann Hypothesis (RH). One striking application was a short proof that at least 2/3 of zeta-zeros are simple zeros, the first result of its type. Over the last 50 years, most work on pair correlation of zeta-zeros has continued to assume RH. Here we show that if RH could be removed from Montgomery’s simple zero proof, then this would also give a proof that 2/3 of the zeros are simple and on the critical line. This idea has been applied in several recent papers to obtain other results on the zeros.
@@ 1. Introduction
In 1973 Montgomery [Mon73] proved, assuming the Riemann Hypothesis (RH), that at least 2/3 of the zeros of the Riemann zeta-function are simple. Prior to this, it was not known either unconditionally or on RH that there were even infinitely many simple zeros. Currently, by a refined version of Levinson’s method, K. Pratt, N. Robles, A. Zaharescu, and D. Zeindler [PRZZ20] in 2020 proved that at least $41.7\%$ of the zeros of $\zeta(s)$ are on the critical line, and at least $40.7\%$ of the zeros are on the critical line and simple. By large scale computation with careful error analysis, Platt and Trudgian [PT21] in 2021 showed that in the critical strip up to height 3,000,175,332,800, there are exactly 12,363,153,437,138 zeros and all of them are on the critical line and are simple. Conditionally, under the RH assumption one can obtain that at least $67.92\%$ of the zeros of $\zeta(s)$ are simple using Montgomery’s pair correlation method [CGdL20], and with a different method [CGG98, BHB13] one obtains at least $70.37\%$ of the zeros are simple.
@@ 2. Zeros of the Riemann zeta-function
since this shows that $(1-2^{1-\sigma})\zeta(\sigma)>0$ so that $\zeta(\sigma)<0$ on the real line for $0<\sigma<1$. We also see immediately from the functional equation that $\zeta(0)=-1/2$.
If an analytic function $F(s)$ is real on a segment of the real axis it will satisfy the reflection principle that $F(\bar{s})=\overline{F(s)}$ in the symmetric region on either side of that segment where $F$ is analytic. Denoting the complex zeros of the Riemann zeta-function by $\rho=\beta+i\gamma$, we see that $\zeta(\rho)=0=\zeta(\bar{\rho})$ and the functional equation gives as well $\zeta(\rho)=0=\zeta(1-\rho)$. Thus, all nontrivial zeros come in pairs symmetric with the real axis, therefore we have the zeros $\rho=\beta+i\gamma$ and $\bar{\rho}=\beta-i\gamma$, and zeros with $\beta\neq 1/2$ come in symmetric quadruples with respect to both the real axis and the critical line (namely, the vertical line $\sigma=1/2$): $\rho=\beta+i\gamma,\ \bar{\rho}=\beta-i\gamma,\ 1-\rho=1-\beta-i\gamma,\ 1-\bar{\rho}=1-\beta+i\gamma$. This elementary fact turns out to be fundamental later in this paper.
3. Counting zeros with multi-sets
Since the zeros are symmetric about the real axis, we only need to study the zeros above the real axis. Analytic function have their zeros counted by the argument principle, which counts zeros with their multiplicity. Thus a zero $\rho$ with multiplicity $m_{\rho}$ is counted as $m_{\rho}$ zeros. Therefore, rather than using “sets” of zeros we will use in this paper multi-sets of zeros where a zero $\rho$ with multiplicity $m_{\rho}$ has $m_{\rho}$ copies of itself in the multiset. Thus we replace the set $Z(T)=\{\rho:\ 0<\gamma\leq T\}$ of complex zeros $\rho$ up to height $T$ by the multiset
(3.1) $\mathcal{Z}(T)=\{\rho:0<\gamma\leq T,\ \rho\ \text{has}\ m_{\rho}\ \text{copies}\}.$
The set $Z(T)$ is called the support of $\mathcal{Z}(T)$, written $\text{Supp}(\mathcal{Z}(T))$, and is the set of distinct zeros.
Let $N(T)$ denote the number of zeros in the critical strip with $0<\gamma\leq T$. In 1905 von Mangoldt proved that, for $T\geq 3$,
(3.2) $N(T)=|\mathcal{Z}(T)|=\sum_{\rho\in\mathcal{Z}(T)}1=\frac{T}{2\pi}\log\frac{T}{2\pi}-\frac{T}{2\pi}+O(\log{T}),$
@@ 4. Montgomery’s RH proof of simple zeros
(4.1) $\sum_{\begin{subarray}{c}\rho,\rho^{\prime}\in\mathcal{Z}(T)\\ \gamma=\gamma^{\prime}\end{subarray}}1\leq\sum_{\rho,\rho^{\prime}\in\mathcal{Z}(T)}\left(\frac{\sin(\frac{1}{2}(\gamma-\gamma^{\prime})\log{T})}{\frac{1}{2}(\gamma-\gamma^{\prime})\log{T}}\right)^{2}=\left(\frac{4}{3}+o(1)\right)\frac{T}{2\pi}\log{T},\qquad\text{as}\quad T\to\infty.$
Here the first inequality is trivially true. The asymptotic formula for the Fejér kernel sum over differences of zeros was proved by Montgomery assuming RH; the interested reader may see the proof in [Mon73]. The Fejér kernel sum itself is obtained without RH from the formula
(4.2) $\sum_{\rho,\rho^{\prime}\in\mathcal{Z}(T)}\left(\frac{\sin(\frac{1}{2}(\gamma-\gamma^{\prime})\log{T})}{\frac{1}{2}(\gamma-\gamma^{\prime})\log{T}}\right)^{2}=2\int_{0}^{1}\bigg|\sum_{\rho\in\mathcal{Z}(T)}T^{i\alpha\gamma}\bigg|^{2}(1-\alpha)\,d\alpha,$
@@
(4.3) $2\,{\rm Re}\int_{0}^{1}e^{iw\alpha}(1-\alpha)\,d\alpha=\left(\frac{\sin(\frac{1}{2}w)}{\frac{1}{2}w}\right)^{2}.$
Thus in (4.1) the double sums are over two copies of the sequence of zeros with $0<\gamma\leq T$, where each sequence already has $m_{\gamma}$ copies of each zero of multiplicity $m_{\gamma}$.
Now consider the first sum in (4.1) with the condition that $\gamma=\gamma^{\prime}$. If we think of the terms as ordered pairs $(\gamma,\gamma^{\prime})$ where $\gamma=\gamma^{\prime}$, then a simple zero gets counted once, but a double zero with $m_{\gamma}=2$ gets counted 4 times, since $\gamma_{1}=\gamma_{2}=\gamma_{1}^{\prime}=\gamma_{2}^{\prime}$ and we get the 4 terms $(\gamma_{1},\gamma_{1}^{\prime})$, $(\gamma_{1},\gamma_{2}^{\prime})$, $(\gamma_{2},\gamma_{1}^{\prime})$, $(\gamma_{2},\gamma_{2}^{\prime})$. Thus a single zero of multiplicity $m_{\gamma}$ gets counted $m_{\gamma}^{2}$ times. The proof of Theorem 1 is now just three lines.
@@
Since by RH, $\gamma=\gamma^{\prime}$ is true if and only if $\rho=\rho^{\prime}$, we have
(4.4) $\sum_{\begin{subarray}{c}\rho,\rho^{\prime}\in\mathcal{Z}(T)\\ \gamma=\gamma^{\prime}\end{subarray}}1=\sum_{\rho\in\mathcal{Z}(T)}m_{\rho}\leq\left(\frac{4}{3}+o(1)\right)\frac{T}{2\pi}\log{T},\qquad\text{as}\qquad T\to\infty.$
Note the simple inequality
(4.5) $\sum_{\begin{subarray}{c}\rho\in\mathcal{Z}(T)\\ \rho\ \text{simple}\end{subarray}}1+\sum_{\rho\in\mathcal{Z}(T)}m_{\rho}=2\sum_{\begin{subarray}{c}\rho\in\mathcal{Z}(T)\\ m_{\rho}=1\end{subarray}}1+\sum_{\begin{subarray}{c}\rho\in\mathcal{Z}(T)\\ m_{\rho}\geq 2\end{subarray}}m_{\rho}\geq 2\sum_{\rho\in\mathcal{Z}(T)}1=2N(T).$
Then using (3.2) and (3.3), we have as $T\to\infty$,
$\begin{split}\sum_{\begin{subarray}{c}\rho\in\mathcal{Z}(T)\\ \rho\ \text{simple}\end{subarray}}1&\geq 2N(T)-\sum_{\rho\in\mathcal{Z}(T)}m_{\rho}\geq\left(2-\frac{4}{3}+o(1)\right)\frac{T}{2\pi}\log{T}=\left(\frac{2}{3}+o(1)\right)\sum_{\rho\in\mathcal{Z}(T)}1.\end{split}$
@@ 5. Montgomery’s Simple Zero Method if RH is not assumed
For 50 years it never seemed to occur to anyone (including the first-named author of this paper) to seriously think about this proof when RH is not assumed. As we will see in a moment, the key idea here is that, despite the title of this section, not assuming RH is not enough; what you have to do is take seriously the possibility that RH is false. The result of this mental block is we spent almost two years proving (4.1) holds with conditions weaker than RH without seeing the connection to zeros on the critical line, which we call for brevity critical zeros in the rest of this exposition.
Recall for the first sum in (4.1) and (4.4) we saw that, assuming RH, the condition $\gamma=\gamma^{\prime}$ is equivalent to $\rho=\rho^{\prime}$. But if RH is false then this is also sometimes false, since then there are in $\mathcal{Z}(T)$ at least two symmetric zeros $\rho=\beta+i\gamma$ and $1-\overline{\rho}=1-\beta+i\gamma$ with $\beta\neq 1/2$. Hence, if $\beta\neq 1/2$ then the pairs of zeros $(\beta+i\gamma,1-\beta+i\gamma)$ and $(1-\beta+i\gamma,\beta+i\gamma)$ are also solutions of the equation $\gamma=\gamma^{\prime}$ on the horizontal line $t=\gamma$. We call these “symmetric diagonal terms”. Thus, in addition to the diagonal terms $\rho=\rho^{\prime}$ in (4.4), which occur for every zero whether RH is assumed or not, we also need to include the symmetric diagonal terms for zeros with $\beta\neq 1/2$. This is not the end. Since we are working with pairs of zeros, if there are three or more zeros on the same horizontal line, we can also have terms which are not diagonal and not symmetric.
To conclude, we have
(5.1) $\sum_{\begin{subarray}{c}\rho,\rho^{\prime}\in\mathcal{Z}(T)\\ \gamma=\gamma^{\prime}\end{subarray}}1\ =\ \sum_{\rho\in\mathcal{Z}(T)}m_{\rho}+\sum_{\begin{subarray}{c}\rho\in\mathcal{Z}(T)\\ \beta\neq\frac{1}{2}\end{subarray}}m_{\rho}+\sum_{\begin{subarray}{c}\rho,\rho^{\prime}\in\mathcal{Z}(T)\\ \ \beta+\beta^{\prime}\neq 1,\,\gamma=\gamma^{\prime}\end{subarray}}1,$
where the first sum on the right-hand side comes from the diagonal terms, the second is given by the symmetric diagonal terms, and the rest in the third sum are non-symmetric horizontal terms.
@@
Suppose there exists a constant $\mathbf{C}$ where $1\leq\mathbf{C}<2$ and that, as $T\to\infty$,
(5.2) $\sum_{\begin{subarray}{c}\rho,\rho^{\prime}\in\mathcal{Z}(T)\\ \gamma=\gamma^{\prime}\end{subarray}}1\leq\Big(\mathbf{C}+o(1)\Big)\frac{T}{2\pi}\log{T}.$
Then asymptotically at least the proportion $2-\mathbf{C}$ of the zeros of $\zeta(s)$ are simple, and at least the proportion $2-\mathbf{C}$ of the zeros of $\zeta(s)$ are on the critical line.
We first remark on the restrictions on the constant $\mathbf{C}$ in our assumption. Note that
(5.3) $\sum_{\begin{subarray}{c}\rho,\rho^{\prime}\in\mathcal{Z}(T)\\ \gamma=\gamma^{\prime}\end{subarray}}1~\geq~\sum_{\begin{subarray}{c}\rho,\rho^{\prime}\in\mathcal{Z}(T)\\ \rho=\rho^{\prime}\end{subarray}}1~=~\sum_{\rho\in\mathcal{Z}(T)}m_{\rho}~\geq~\sum_{\rho\in\mathcal{Z}(T)}1=N(T)~\sim~\frac{T}{2\pi}\log{T}$
as $T\to\infty$. Thus $\mathbf{C}$ in our assumption has to satisfy $\mathbf{C}\geq 1$. We also restrict $\mathbf{C}<2$, since for $\mathbf{C}\geq 2$ the method fails to give a positive proportion.
@@
(5.4) $\sum_{\rho\in\mathcal{Z}(T)}m_{\rho}+\sum_{\begin{subarray}{c}\rho\in\mathcal{Z}(T)\\ \beta\neq\frac{1}{2}\end{subarray}}m_{\rho}+\sum_{\begin{subarray}{c}\rho,\rho^{\prime}\in\mathcal{Z}(T)\\ \ \beta+\beta^{\prime}\neq 1,\,\gamma=\gamma^{\prime}\end{subarray}}1\leq\Big(\mathbf{C}+o(1)\Big)\frac{T}{2\pi}\log{T},\qquad\text{as}\quad T\to\infty.$
@@
$\sum_{\rho\in\mathcal{Z}(T)}m_{\rho}\leq\Big(\mathbf{C}+o(1)\Big)\frac{T}{2\pi}\log{T},\qquad\text{as}\quad T\to\infty,$
which is (4.4) with $\mathbf{C}=4/3$, and the identical argument (4.5) used there gives
(5.5) $\sum_{\begin{subarray}{c}\rho\in\mathcal{Z}(T)\\ \rho\ \text{simple}\end{subarray}}1\geq\Big(2-\mathbf{C}+o(1)\Big)\frac{T}{2\pi}\log{T},\qquad\text{as}\quad T\to\infty,$
which proves that the proportion of simple zeros of $\zeta(s)$ is $\geq 2-\mathbf{C}$.
Next, for the symmetric diagonal terms, we have, as $T\to\infty$,
$\sum_{\begin{subarray}{c}\rho\in\mathcal{Z}(T)\\ \beta\neq 1/2\end{subarray}}m_{\rho}\leq\Big(\mathbf{C}+o(1)\Big)\frac{T}{2\pi}\log{T}-\sum_{\rho\in\mathcal{Z}(T)}m_{\rho}.$
Recalling the lower bound in (5.3), we obtain
(5.6) $\sum_{\begin{subarray}{c}\rho\in\mathcal{Z}(T)\\ \beta\neq 1/2\end{subarray}}m_{\rho}\leq\Big(\mathbf{C}-1+o(1)\Big)\frac{T}{2\pi}\log{T},\qquad\text{as}\quad T\to\infty.$
@@
$\sum_{\begin{subarray}{c}\rho\in\mathcal{Z}(T)\\ \beta=1/2\end{subarray}}1=\sum_{\rho\in\mathcal{Z}(T)}1-\sum_{\begin{subarray}{c}\rho\in\mathcal{Z}(T)\\ \beta\neq 1/2\end{subarray}}1\geq N(T)-\sum_{\begin{subarray}{c}\rho\in\mathcal{Z}(T)\\ \beta\neq 1/2\end{subarray}}m_{\rho}\geq\Big(2-\mathbf{C}+o(1)\Big)\frac{T}{2\pi}\log{T}$
@@
as $T\to\infty$, and we have proved that the proportion of zeros of $\zeta(s)$ on the critical line is $\geq 2-\mathbf{C}$. ∎
6. Simple and critical zeros using diagonal and symmetric diagonal terms
While we cannot improve either result in Theorem 2, we can obtain further results on zeros that are both simple and critical, simple or critical, and also the average of the proportion of simple zeros and the proportion of critical zeros.
Theorem 3.
Suppose there exists a constant $\mathbf{C}\geq 1$ and that, as $T\to\infty$,
(6.1) $\sum_{\rho\in\mathcal{Z}(T)}m_{\rho}+\sum_{\begin{subarray}{c}\rho\in\mathcal{Z}(T)\\ \beta\neq\frac{1}{2}\end{subarray}}m_{\rho}\leq\Big(\mathbf{C}+o(1)\Big)\frac{T}{2\pi}\log{T}.$
Then asymptotically,
  1. (i)